# Hotfixing Large Language Models for Code: How Far Can Parameter-Efficient Fine-Tuning Go?

### v2

@@ Abstract.
Large Language Models for Code (LLM4Code) have become an integral part of developers’ workflows, assisting with tasks such as code completion and generation. However, these models are found to exhibit undesired behaviors after their release, like generating buggy code, due to their extensive training on vast amounts of source code that contain such buggy code. The training data (usually coming from open-source software) keeps evolving, e.g., developers fix the buggy code. However, adapting such evolution to mitigate LLM4Code’s undesired behaviors is non-trivial, as retraining models on the updated dataset usually takes much time and resources. This motivates us to propose the concept of hotfixing LLM4Code, mitigating LLM4Code’s undesired behaviors effectively and efficiently with minimal negative effects.
This paper mainly focuses on hotfixing LLM4Code to make them generate less buggy code and more fixed code. We begin by demonstrating that models from the popular CodeGen family frequently generate buggy code. Then, we define three learning objectives in hotfixing and design multiple loss functions for each objective: (1) learn the desired behaviors, (2) unlearn the undesired behaviors, and (3) retain knowledge of other code. We evaluate four different fine-tuning techniques for hotfixing the models and gain the following insights. Optimizing these three learning goals together, using LoRA (low-rank adaptation), effectively influences the model’s behavior. Specifically, it increases the generation of fixed code by up to 108.42% and decreases the generation of buggy code by up to 50.47%. Statistical tests confirm that hotfixing does not significantly affect the models’ functional correctness on the HumanEval benchmark. We also show that hotfixing demonstrates strong time efficiency.
@@ 1. Introduction
Large Language Models for Code (LLM4Code) (Fan et al., 2023; Hou et al., 2023; Zheng et al., 2023) have demonstrated outstanding performance on tasks that span a variety of activities in software engineering, including requirement gathering (Bencheikh and Höglund, 2023), architecture design (Ahmad et al., 2023), code generation (Nijkamp et al., 2023), software testing (Xia and Zhang, 2023), defect prediction (Zhou et al., 2023), stack overflow post analysis (He et al., 2022), etc. Many tools (e.g., GitHub Copilot (git, [n. d.]), AWS CodeWhisperer (cod, [n. d.]), etc), which can complete and generate code, have been integrated into widely-used IDEs such as Visual Studio Code to assist developers in their daily tasks.
Despite their strong performance, LLM4Code are usually found to exhibit undesired behaviors. For example, Jesse et al. (Jesse et al., 2023) find that LLM4Code can memorize and produce many known simple, stupid bugs (i.e., bugs that can be fixed with single-line changes, also known as SStuBs (Karampatsis and Sutton, 2020)). As another example, researchers have found that LLM4Code can generate a large number of privacy-sensitive information (Yang et al., 2024b), like email addresses and API keys. Considering the large population of users of LLM4Code, [^1] Reports show that GitHub Copilot has already been used by over 1.2 million developers. there exists a key requirement to update the model as soon as possible to minimize the potential negative impacts on more users caused by these undesired behaviors. One intuitive way is to retrain the model on the updated dataset, by adding the fixed code or replacing the sensitive information with placeholders in the training data, which however, is extremely time-consuming and computationally expensive.
@@
This practical requirement motivates us to explore the concept of “hotfixing” LLM4Code. Hotfix, in the context of traditional software engineering, is an unplanned improvement to remediate the unwanted symptoms of the critical issue quickly (Hanna et al., 2024). Hotfix emphasizes less on correctness under all conditions but more on the time required to generate a plausible patch that hides the critical symptom without breaking the system. In the context of LLM4Code, retraining the model thoroughly can be quite time-consuming, we hope to first hotfix the model to mitigate the undesired behaviors and thus leave time for a complete fix (retraining). A good hotfix for LLM4Code should satisfy the following properties: (1) efficiency: the time and effort to obtain the hotfix should be significantly faster than retraining the model; (2) effectiveness: the hotfix should effectively mitigate the behavior while causing minimal impact to the functional correctness; and (3) easy deployment: the hotfix should minimize disruption and complexity during deployment.
He and Vechev (He and Vechev, 2023) propose the pioneering study on updating LLM4Code to reduce the generation of insecure code. They define three targets for such updates: the updated model should (1) produce more secure code, (2) produce less insecure code, and (3) remain the same for other code. To achieve these targets, they design a combination of three loss functions to train contrastive prefix to update the models. This paper extends their study by implementing hotfixing to mitigate another important type of undesired behavior: LLM4Code can complete many known buggy code; we aim at making LLM4Code not complete buggy code but complete the corresponding fixed code. First, we replicate the experiments of Jesse et al. (Jesse et al., 2023), confirming that models from a popular family (CodeGen) can memorize and produce many bugs, which is treated as undesired behaviors that should be mitigated. These known bugs have been fixed in open-source software; thus we hotfix the models and steer them to generate the corresponding fixed versions. To achieve this goal, we construct a dataset consisting of pairs of two code snippets: one containing bugs and the other with corresponding bugs being fixed. Then, we compute the diff (to get the deleted buggy code and the added fixed code) of each pair.
We experiment a set of parameter-efficient fine-tuning (PEFT) methods to hotfix the models, which are designed to update the model with a small number of parameters, resulting in efficiency gains and making the deployment easier. We investigate a family of popular LLM4Code, CodeGen (Nijkamp et al., 2023), which is widely adopted in multiple studies (Jesse et al., 2023; Yang et al., 2024b; Li et al., 2023; Allal et al., 2023). By conducting experiments on models of two sizes (CodeGen-350M and CodeGen-2B), we gain the following insights about hotfixing LLM4Code. The optimal hotfixing setting (combining all three objectives and using LoRA for fine-tuning) can increase the generation of fixed code by up to 108.42% and decrease the generation of buggy code by up to 50.47%, which is more effective than the mitigation strategy adopted by Jesse et al. (Jesse et al., 2023). Although hotfixing reduces the model’s functional correctness on the HumanEval benchmark, statistical tests show that the difference is not statistically significant ($p$-values are greater than 0.05).
@@
  • Task: We propose the novel concept of hotfixing LLM4Code, aiming at mitigating the LLM4Code’s undesired behaviors effectively and efficiently. This is a new task in LLM4Code maintenance.
  • Effectiveness: We implement a hotfix strategy by extending the work by He and Vechev (He and Vechev, 2023) with more PEFT methods on a new task, showing that it can effectively mitigate undesired behaviors with no significant impact on the functional performance.
  • Efficiency: Hotfixing LLM4Code demonstrates efficiency. Specifically, it takes 5 minutes to hotfix CodeGen-350M on a single GPU.
  • Application: We discuss the potential application of hotfixing in more scenarios, like deprecated API updates and customized code generation, as well as deployment in practice.
Paper Structure. Section 2 provides the background information of this study. We explain our methodology in Section 3, including motivating scenarios and learning objectives. Section 4 documents the experiment settings, and we present our results in Section 5. We conduct a case study about reducing privacy leakage in Section LABEL:sec:case-study. After presenting additional discussion in Section 6, we explain related works in Section 7. Section 8 concludes the findings of our paper, and outlines future studies.
@@ 2.1. LLM-based Code Generators
According to a recent survey of LLM4Code (Hou et al., 2023), code generation models are largely decoder-only models (Nijkamp et al., 2023; Fried et al., 2023; Li et al., 2023), which are typically based on the Transformer architecture (Vaswani et al., 2017). These models are pre-trained on a large corpus of datasets in an unsupervised manner to learn a probability distribution that predicts the likelihood of the next token, given the context (also known as prompt). Formally speaking, a model takes in a series of tokens represented as $x=(x_{1},x_{2},\cdots,x_{t})$. It then produces a probability distribution $P(x_{t+1}|x_{1},x_{2},\cdots,x_{t})$, which estimates the likelihood of the next token in the sequence being $x_{t+1}$. The computation of the probability distribution can be factorized into two steps: (1) a function $f_{h}$ that computes the hidden states of the input sequence and (2) a function $f_{p}$ that computes the probability distribution of the next token based on the hidden states.
@@
The models produce subsequent tokens in an autoregressive manner. The model first predicts the probability distribution of the next token, and then selects the token with the highest probability as the next token $x_{t+1}$. The new token is then fed into the model to predict the next token $x_{t+2}$. The process is repeated until the model predicts a special token (e.g., the end of sequence token) that indicates the end of the sequence or other stopping criteria (e.g., the maximal number of generated tokens) are met.
@@ 2.2. Parameter-Efficient Fine-Tuning
To make the hotfix efficient and easy to deploy, this paper employs parameter-efficient fine-tuning (PEFT) techniques to hotfix LLM4Code. PEFT methods train LLM4Code by only updating a small number of parameters rather than updating the entire model, thus reducing the computational cost and time required for training. Here the updated parameters are the hotfix to mitigate models’ undesired behaviors. One of the representative PEFT method is LOw-Rank Adaptation (LoRA) (Hu et al., 2022a). LoRA freezes the model weights and adds low-rank trainable matrices into the attention layers of Transformer models. By only updating the newly added matrices, LoRA significantly reduces the number of parameters to be trained. IA3 (Liu et al., 2022) aims to improve LoRA and further reduces the amount of trainable parameters. Another PEFT method is prefix-tuning (Li and Liang, 2021), which trains a set of virtual tokens added to the input tokens of the LLM. Dettmers et al. (Dettmers et al., 2023) combine LoRA with quantization, leading to less GPU memory consumption when training LoRA matrices.
Following a recent study (Weyssow et al., 2024) on evaluating PEFT methods on LLM4Code, we choose LoRA, IA3, prefix-tuning, and QLoRA as the PEFT methods to obtain hotfix LLM4Code in this paper.
@@ 3. Methodology
In this section, we begin by describing a motivating scenario where code LLM4Code produce undesired code. We discuss how to collect the dataset and evaluate such behaviors. We then explain the learning objectives and the methods to update the models.
Figure 1. An overview of Hotfixing LLM4Code. The hotfix is obtained by using parameter-efficient fine-tuning (PEFT) methods to specifically learn the difference between the desired and undesired behaviors. Depending on the PEFT methods, the hotfix can be in different forms, e.g., LoRA matrix. The hotfix is much smaller than the original model. The hotfix using LoRA is only 13.9MB for CodeGen-2B.
@@ 3.1. Motivating Scenario
Researchers find that there exists much buggy code in the open-source repositories, which is then collected to appear in LLM4Code training data. The models can learn from the training data to generate buggy code. For example, Jesse et al. (Jesse et al., 2023) investigate a specific category of bugs, namely, simple, stupid bugs (SStuBs) (Karampatsis and Sutton, 2020), and find that code generators (like CodeGen (Nijkamp et al., 2023)) can generate known SStuBs, and at a high rate (around twice as often as they produce known correct, bug-fixing code). Listing 1 shows an example of buggy code generated by CodeGen-350M.
The code snippet before the part highlighted in red is the prompt sent to the CodeGen-350M model, and the part highlighted in red (i.e., stacktrace.indexOf(’:’)) is the generated code, which is buggy. [^3] https://github.com/ACRA/acra Although the bug has been fixed in the original repository (the fixed code is firstLine.indexOf(’:’)), the model still generates the buggy version. As shown in the upper half of Figure 1, a model is trained on a dataset containing buggy code. To alleviate this issue of memorizing known bugs, one option is to replace the buggy code with the fixed code in the training data and retrain the model, which can be considered as a ‘complete’ fix. However, retraining is computationally expensive and time-consuming. Xu et al. (Xu et al., 2022) report that it takes 6 weeks to train a 2.7B parameters PolyCoder using 8 Nvidia RTX 8000 GPUs. Besides, the ‘other code’ part in Figure 1, i.e., the remaining part of the original training data, may not be available. For example, the training data of two models investigated in this paper is not publicly available.
In this paper, we learn the ‘hotfix’ by guiding the model to learn the differences between the buggy code and the fixed code. On the one side, we only need a small amount of changed code (explained in Section 3.2) to guide the model to learn different objectives (explained in Section 3.3). On the other side, we explore different parameter-efficient fine-tuning methods to update the model (explained in Section 3.4).

### v4

@@ Abstract.
Large Language Models for Code (LLM4Code) have become indispensable tools for developers, aiding in tasks like code generation, completion, and analysis. However, these models face a significant challenge: adapting to the ever-changing software. Critical updates, such as bug fixes, must be promptly reflected in LLM4Code to ensure they generate correct and privacy-preserving code. Adapting to these updates and requirements by retraining LLM4Code from scratch is impractical due to the immense time and resources required.
To address this, we propose the concept of hotfixing LLM4Code — a model maintenance activity that implements time-critical improvements efficiently and effectively, targeting specific issues without extensive retraining. We explore both online hotfixing, which modifies model behavior through prompt adjustments without interrupting service, and offline hotfixing, which updates model parameters using parameter-efficient fine-tuning techniques to specifically guide models to learn the updates. Our experiments focus on hotfixing LLM4Code to (1) generate fixed code instead of outdated, buggy code, and (2) produce privacy-preserving code completions by replacing sensitive information with placeholders. Evaluating on the CodeGen family of models, we extend existing strategies with more fine-tuning methods like LoRA. Our results show that hotfixing increases the generation of fixed code by up to 108.42% and decreases buggy code generation by up to 50.47%. Additionally, it reduces the leakage of email addresses by up to 99.30%, with the process taking as little as 5 minutes on a single GPU. These findings demonstrate that hotfixing is an effective and efficient solution for updating LLM4Code, enabling them to keep pace with software evolution while minimizing disruption and resource consumption.
@@
ccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software notations and tools
@@ 1. Introduction
“The only thing that never changes is that everything changes.”
Louis L’Amour
@@
Software never stops changing: requirements evolve, third-party libraries update, new programming languages emerge, design patterns change, coding standards differ, and licenses vary. Software and developers need to adapt to this ever-changing environment in a coordinated manner. Now a new entity, Large Language Models for Code (LLM4Code) (Hou et al., 2024), has been introduced to enhance the synergy between software and developers. LLM4Code consumes vast resources and time to learn from massive datasets and provides developers with various forms of assistance, such as requirement gathering (Bencheikh and Höglund, 2023), architecture design (Ahmad et al., 2023), code generation (Nijkamp et al., 2023), software testing (Xia and Zhang, 2023), defect prediction (Zhou et al., 2023), stack overflow post analysis (He et al., 2022), etc. Many tools (e.g., GitHub Copilot (git, [n. d.]), AWS CodeWhisperer (cod, [n. d.]), etc.), which can complete and generate code, have been integrated into widely-used IDEs such as Visual Studio Code to support developers in their daily tasks.
It leads to a straightforward challenge: how to adapt LLM4Code to the ever-changing software environment? Some critical changes in software should be reflected in the models. For example, Jesse et al. (Jesse et al., 2023) find that mainstream LLM4Code models (like CodeGen (Nijkamp et al., 2023) and Codex (Chen et al., 2021)) often output buggy code verbatim, even though these bugs have been fixed. Since these models are trained on data collected before the fixes, they do not reflect the updates. Another concern is that LLMs may memorize private information from their training data. For example, researchers have recently found that LLM4Code can produce privacy-sensitive information in their outputs (Yang et al., 2024c; Niu et al., 2023), like email addresses and API keys. Although this information might still exist in the software, outputting it can violate privacy regulations like General Data Protection Regulation (GDPR) and California Consumer Privacy Act (CCPA), potentially harming the model provider’s reputation. This motivates updating models to produce expected behaviors—such as producing fixed rather than buggy code and replacing emails with placeholders to generate privacy-preserving completions.
However, updating LLM4Code to adapt to changes in the software environment is non-trivial. Retraining these models can be extremely time-consuming and resource-intensive, making it difficult to keep up with the rapid pace of software evolution. It is reported that, on average, 10 million contributions are made to GitHub every day. [^1] The contributions include commits, issues, pull requests, discussions, gists, pushes, and pull request reviews. Data is fetched from https://octoverse.github.com/2022/developer-community Additionally, undesired behaviors in LLM4Code are constantly being discovered. For example, researchers have identified various flaws in LLMs, such as generating buggy code (Jesse et al., 2023), exposing private information (Yang et al., 2024c; Niu et al., 2023), and misusing APIs. Considering the large population of users of LLM4Code, [^2] Reports show that GitHub Copilot has already been used by over 1.2 million developers. developers to update LLM4Code as soon as possible to minimize the potential negative impacts caused by these undesired behaviors. This means that the request to update LLM4Code can be frequent and can address various issues, which highlights the requirements for efficient and effective updates.
@@
This practical requirement motivates us to explore the novel concept of “hotfixing” LLM4Code. A hotfix, in the context of traditional software engineering, is an unplanned improvement to quickly remediate the unwanted symptoms of a critical issue (Hanna et al., 2024). A hotfix emphasizes less on correctness under all conditions but more on the time required to generate a plausible patch that hides the critical symptom without breaking the system. Following existing studies (Gupta et al., 2008; Agarwal and Garg, 2014; Hanna et al., 2024; Hanna and Petke, 2024), we define hotfixing LLM4Code as follows:
Definition: Hotfixing Large Language Model for Code (LLM4Code) Hotfixing LLM4Code is a model maintenance activity to implement a time-critical improvement that is temporary, small in size, fast to obtain, and targets a specific issue in LLM4Code.
Existing hotfixing techniques (Salehi and Pattabiraman, 2022; Novark et al., 2008; Zhang and Yin, 2014) for conventional software, where the logic is defined using programming languages, are usually not applicable to LLM4Code, which is defined by millions and even trillions of parameters. Thus, we conduct a series of comprehensive experiments to explore appropriate strategies for hotfixing LLM4Code. Researchers have proposed various strategies to alter the behavior of LLM4Code, which we can leverage to hotfix the models. Depending on how the hotfix is deployed into the model to take effect, we divide the hotfix into two categories: online hotfix and offline hotfix. The online hotfix means that the hotfix is deployed into the model without interrupting the model’s service. As the behavior of LLM4Code is affected by the prompts, the online hotfix can be achieved by modifying the prompts sent to the model. For example, Jesse et al. (Jesse et al., 2023) add natural language describing the desired completion to guide the model to generate the fixed code. Huang et al. (Huang et al., 2023a) alter the prompts to mitigate the bias in LLM4Code’s outputs. The offline hotfix requires interrupting and restarting the model to take effect, e.g., updating the model parameters and reloading the model. For example, He and Vechev (He and Vechev, 2023) use prefix-tuning (Li and Liang, 2021) to drive the model to generate more secure code and less vulnerable code.
Our experiments focus on hotfixing LLM4Code to adapt to two critical changes: (1) learning from the updated dataset to generate fixed code instead of buggy code, and (2) learning to generate privacy-preserving completions by automatically localizing and replacing sensitive information with placeholders. We investigate a family of popular LLM4Code, CodeGen (Nijkamp et al., 2023), which is widely adopted in multiple studies (Jesse et al., 2023; Yang et al., 2024c; Li et al., 2023; Allal et al., 2023). Our experiment includes CodeGen models of two sizes (CodeGen-350M and CodeGen-2B). We use the strategies employed by Jesse et al. (Jesse et al., 2023) and He and Vechev (He and Vechev, 2023) as representatives for online hotfix and offline hotfix, respectively. We also extend the method by He and Vechev (He and Vechev, 2023) by exploring more parameter-efficient fine-tuning (PEFT) methods and analyzing how different learning objectives can affect the hotfixing results. We find that hotfixing using LoRA (Hu et al., 2021) can increase the generation of fixed code by up to 108.42% and decrease the generation of buggy code by up to 50.47%, which is more effective than the online mitigation strategy adopted by Jesse et al. (Jesse et al., 2023). We also find that the email leakage of CodeGen-350M reduces by up to 99.30% after hotfixing, which only takes 5 minutes on a single GPU with 24 GB memory.
@@
  • Task: We propose the novel concept of hotfixing LLM4Code, a model maintenance activity to implement a time-critical improvement that is temporary, small in size, fast to obtain, and targets a specific issue in LLM4Code.
  • Effectiveness: We evaluate both online and offline hotfixing by extending the existing strategies used by Jesse et al. (Jesse et al., 2023) and He and Vechev (He and Vechev, 2023), showing that LLM4Code can be hotfixed effectively to mitigate undesired behaviors with no significant impact on functional performance.
  • Efficiency: We evaluate the efficiency of hotfixing LLM4Code. Specifically, it takes 5 minutes to hotfix CodeGen-350M on a single GPU.
  • Application: We discuss the potential application of hotfixing in more scenarios, such as deprecated API updates and customized code generation, as well as deployment in practice.
Paper Structure. Section 2 provides the background for this study. We explain our methodology in Section 3. We present our results in Section 4. We conduct a case study on reducing privacy leakage in Section 5. After presenting additional discussion in Section 6, we explain related works in Section 7. Section 8 concludes the findings of our paper and outlines future studies.
@@ 2.1. LLM-based Code Generators
According to a recent survey of LLM4Code (Hou et al., 2024), code generation models are largely decoder-only models (Nijkamp et al., 2023; Fried et al., 2023; Li et al., 2023), which are typically based on the Transformer architecture (Vaswani et al., 2017). These models are pre-trained on a large corpus of datasets in an unsupervised manner to learn a probability distribution that predicts the likelihood of the next token, given the context (also known as a prompt). Formally speaking, a model takes in a series of tokens represented as $x=(x_{1},x_{2},\cdots,x_{t})$. It then produces a probability distribution $P(x_{t+1}|x_{1},x_{2},\cdots,x_{t})$, which estimates the likelihood of the next token in the sequence being $x_{t+1}$. The computation of the probability distribution can be factorized into two steps: (1) a function $f_{h}$ that computes the hidden states of the input sequence and (2) a function $f_{p}$ that computes the probability distribution of the next token based on the hidden states.
@@
The models produce subsequent tokens in an autoregressive manner. The model first predicts the probability distribution of the next token and then selects the token with the highest probability as the next token $x_{t+1}$. The new token is then fed into the model to predict the next token $x_{t+2}$. The process is repeated until the model predicts a special token (e.g., the end of sequence token) that indicates the end of the sequence or other stopping criteria (e.g., the maximum number of generated tokens) are met.
@@ 2.2. Parameter-Efficient Fine-Tuning
To make the hotfix efficient and easy to deploy, this paper employs parameter-efficient fine-tuning (PEFT) techniques to hotfix LLM4Code. PEFT methods train LLM4Code by only updating a small number of parameters rather than updating the entire model, thus reducing the computational cost and time required for training. Here, the updated parameters are the hotfix to mitigate the model’s undesired behaviors. One of the representative PEFT methods is Low-Rank Adaptation (LoRA) (Hu et al., 2021). LoRA freezes the model weights and adds low-rank trainable matrices into the attention layers of Transformer models. By only updating the newly added matrices, LoRA significantly reduces the number of parameters to be trained. IA3 (Liu et al., 2022) aims to improve LoRA and further reduces the amount of trainable parameters. Another PEFT method is prefix-tuning (Li and Liang, 2021), which trains a set of virtual tokens added to the input tokens of the LLM. Dettmers et al. (Dettmers et al., 2023) combine LoRA with quantization, leading to less GPU memory consumption when training LoRA matrices.
Following a recent study (Weyssow et al., 2024) on evaluating PEFT methods on LLM4Code, we chose LoRA, IA3, prefix-tuning, and QLoRA as the PEFT methods to obtain hotfix LLM4Code in this paper.
@@ 3. Methodology
In this section, we begin by describing a motivating scenario where LLM4Code exhibits undesired behaviors and requires hotfixing to adapt to changes in the training data. We discuss how to collect the dataset and evaluate such behaviors. We then explain the methods to hotfix LLM4Code.
Figure 1. An overview of Hotfixing LLM4Code. The hotfix is obtained by using parameter-efficient fine-tuning (PEFT) methods to specifically learn the difference between the desired and undesired behaviors. Depending on the PEFT methods, the hotfix can be in different forms, e.g., a LoRA matrix. The hotfix is much smaller than the original model. The hotfix using LoRA is only 13.9 MB for CodeGen-2B.
@@ 3.1. Motivating Scenario
Researchers find that there is much buggy code in open-source repositories, which is then collected to appear in LLM4Code training data. The models can learn from the training data to generate buggy code. For example, Jesse et al. (Jesse et al., 2023) investigate a specific category of bugs, namely, simple, stupid bugs (SStuBs) (Karampatsis and Sutton, 2020), and find that code generators (like CodeGen (Nijkamp et al., 2023)) can generate known SStuBs at a high rate (around twice as often as they produce known correct, bug-fixing code).
Listing 1 shows an example of buggy code generated by CodeGen-350M. The code snippet before the part highlighted in red is the prompt sent to the CodeGen-350M model, and the part highlighted in red (i.e., stacktrace.indexOf(’:’)) is the generated code, which is buggy. [^4] https://github.com/ACRA/acra Although the bug has been fixed in the original repository (the fixed code is firstLine.indexOf(’:’)), the model is trained on the data before the bugs are fixed and thus still generates the buggy version. As shown in the upper half of Figure 1, a model is trained on a dataset containing buggy code. To alleviate this issue of memorizing known bugs, one option is to replace the buggy code with the fixed code in the training data and retrain the model, which can be considered a ’complete’ fix. However, retraining is computationally expensive and time-consuming. Xu et al. (Xu et al., 2022) report that it takes 6 weeks to train a 2.7B parameter PolyCoder using 8 Nvidia RTX 8000 GPUs. Besides, the ’other code’ part in Figure 1, i.e., the remaining part of the original training data, may not be available. For example, the training data of CodeGen (Nijkamp et al., 2023) is not publicly available. This motivates us to explore how to hotfix LLM4Code to learn the changes in the training data to mitigate the undesired behaviors.
@@
3.2. Online Hotfix Using Prompts
The online hotfix means that the hotfix is deployed into the model without interrupting the model’s service. The behavior of LLM4Code is affected by the prompts, and good prompts demonstrate effectiveness in guiding LLM4Code to produce outputs that are desired by the users. Thus, an online hotfix can be achieved by modifying the prompts sent to the model, which does not require stopping and restarting the model. For this specific task to generate the fixed code, we adopt the strategy used by Jesse et al. (Jesse et al., 2023) (who exposed this undesired behavior) as the representative for the online hotfix. To encourage the models to generate fixed code, Jesse et al. (Jesse et al., 2023) add a natural language comment before the location to be completed to pre-condition the model. They use a CodeTrans model (Elnaggar et al., 2021) fine-tuned on the DeepCom dataset (Hu et al., 2018) as the code summarization tool. CodeTrans is used to generate comments for both the buggy statement and its corresponding fix, which are called BugComment and FixComment, respectively. The experiment results (Jesse et al., 2023) show that FixComment can condition the model to generate more fixed code. For example, after adding FixComment, CodeGen-350M (Nijkamp et al., 2023) generates 7.58% less buggy code and 13.12% more fixed code. This is analogous to the situation where users know what they want to generate and describe the code in natural language to prompt LLM4Code. Jesse et al. (Jesse et al., 2023) make the prompts with FixComment available to the public, which are used in this paper to evaluate the online hotfix.
3.3. Offline Hotfix
The offline hotfix requires interrupting and restarting the model to take effect, which usually involves updating the model parameters and reloading the model. One intuitive strategy to update LLM4Code is to fine-tune LLM4Code on the updated dataset. Recall how LLM4Code generates outputs: given a sequence of tokens as input, the model generates a distribution over the next token. Given a training example input $x$ and its target output $y$, the model computes the probability of outputting $y$ given $x$, denoted by $P(y|x)$. The model learns this training example by maximizing this probability. In practice, this process is conducted by minimizing the negative log-likelihood (NLL) loss function, which is defined as:
(2) $L_{vanilla}=-\frac{1}{N}\sum_{i=1}^{N}\log P(y_{i}|x_{i})$
$N$ is the number of training samples, $x_{i}$ is the input for the $i$-th sample, and $y_{i}$ is the target output for the $i$-th sample. By minimizing this loss function, the model adjusts its parameters to increase the probabilities of the target words, improving its ability to predict the next word in similar sentences. This loss function is commonly used to tune LLM4Code (Weyssow et al., 2024). We call it the Vanilla loss function.
He and Vechev (He and Vechev, 2023) guide the model to assign different importance to different tokens in a training example. To guide the model to generate more secure code, they let the model train on the code snippets that are viewed as secure. In hotfixing, not every token in the entire updated examples is equally important. For example, the target of hotfixing is to generate the fixed code. The newly added fixed code is considered more important than other parts of the not updated code and deserves more attention. We apply this idea to hotfixing LLM4Code to guide the model to learn the added fixed code. The following loss function (He and Vechev, 2023) is used:
(3) $L_{guided}=-\frac{1}{N}\sum w_{t}^{+}\log P(x_{t}|x_{<t})$
In the above equation, the loss value for a token $x_{t}$ is multiplied by a weight $w_{t}^{+}$. This weight indicates whether a token is the newly added code or not, where $w_{t}^{+}=1$ if $x_{t}$ is the newly added code (i.e., the fixed code we want to generate) and $w_{t}^{+}=0$ otherwise. In this way, the loss on other parts of a training example is not used to update the model parameters, and we guide the model to learn the fixed code. We call this loss setting the Guided loss.
Software changes are usually in the format of diffs, which contain both the added code (e.g., fixed code) and the removed code (e.g., buggy code). Thus, to hotfix LLM4Code not only generates the fixed code but also ‘unlearns’ the old buggy code. Similarly to Equation (3), the log-likelihood loss function on the undesired code can be computed as follows:
(4) $L_{unlearn}=-\frac{1}{N}\sum w_{t}^{-}\log P(x_{t}|x_{<t})$
In the above equation, $w_{t}^{-}$ is the weight in computing the loss on token $x_{t}$, where $w_{t}^{-}=1$ if $x_{t}$ is the deleted code and $w_{t}^{-}=0$ otherwise. Unlike the desired code, we want to forget the information of the undesired code, i.e., increase the model loss on the undesired code to avoid it from being generated. However, simply maximizing the loss on the undesired code is ineffective. Our experiment shows that maximizing this loss will make the model lose its ability to generate meaningful code (the hotfixed models generate zero buggy code and zero fixed code), which is not desired in hotfixing. Instead of directly maximizing this loss, He and Vechev (He and Vechev, 2023) minimize the loss function $\frac{L_{guided}}{L_{guided}+L_{unlearn}}$. Intuitively, the optimizer finds a balance between minimizing the loss on desired code $L_{guided}$ and maximizing the loss on undesirable code $L_{unlearn}$. Ideally, to minimize this loss, the optimizer works to reduce $L_{guided}$, which encourages fixed code generation, and to increase $L_{unlearn}$, which discourages the generation of buggy code. To make the model learn the desired code and avoid the undesired code, we define the Dual loss as follows:
(5) $L_{dual}=(L_{guided}+\frac{L_{guided}}{L_{guided}+L_{unlearn}})/2$
Although the loss functions above do not touch the code that is not changed, training may affect the knowledge of this part, which may introduce other undesired behaviors (e.g., harming the correctness of generated code). As an analogy, a commit fixing a bug may introduce another. Although hotfixes put less emphasis on side effects, we expect to avoid such a situation for hotfixes.
The fact that tuning language models on a new dataset can lead to performance degradation in other aspects has been observed in previous studies. For example, Liu et al. (Liu et al., 2016) point out that optimizing models on a new dataset can improve the average performance (e.g., BLEU score) but forgo coherence and fluency. This degeneration is often diagnosed as an effect of deviating too much from the original pre-trained model during optimization. Consequently, we want the model to behave similarly to the original model on the code that is not relevant to the undesired behavior. This is similar to the target of knowledge distillation (Tang et al., 2024; Shi et al., 2024), which aims to train a student model to mimic the behavior of the teacher model. In knowledge distillation, Kullback-Leibler (KL) divergence (Kullback and Leibler, 1951) is widely used to measure the behavioral difference between the two models, and the student model learns to minimize the difference (Shi et al., 2023; Kim and Rush, 2016; Kim et al., 2024). He and Vechev also use it to mitigate the potential functional correctness decrease. The KL divergence loss is defined as follows:
(6) $L_{KL}=\sum D_{KL}(P(x_{t}|x_{<t})||P_{0}(x_{t}|x_{<t}))$
In the above equation, $P_{0}$ is the probability generated by the original model. In Equation (6), the notation $D_{KL}$ represents KL divergence between the original model’s token distribution and the hotfixed model’s token distribution. KL divergence is a measure of how one probability distribution diverges from a second, expected probability distribution. During the hotfixing process, we monitor the token distribution of the original model and the model under hotfixing. We then compute the average KL divergence of each unchanged position in a training example, which is denoted as $L_{KL}$. In our experiment, we also add $L_{KL}$ to each of the loss functions designed above, denoted by Vanilla+KL, Guided+KL, and Dual+KL. For example, the loss function of Guided+KL is $(L_{guided}+L_{KL})/2$ and the loss function of Dual+KL is $(L_{guided}+\frac{L_{guided}}{L_{guided}+L_{unlearn}}+L_{KL})/3$.